# FlowVLA: Thinking in Motion with a Visual Chain of Thought

### v1

@@ Abstract
Many Vision-Language-Action (VLA) models rely on an internal world model trained via next-frame prediction. This approach, however, struggles with physical reasoning as it entangles static appearance with dynamic motion, often resulting in implausible visual forecasts and inefficient policy learning. To address these limitations, we introduce the Visual Chain of Thought (Visual CoT): a pre-training framework that encourages a model to reason about how a scene evolves before predicting what it will look like. We instantiate this principle in FlowVLA, which predicts a future frame ($v_{t+1}$) only after generating an intermediate optical flow representation ($f_{t}$) that encodes motion dynamics. This $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$” reasoning process is implemented within a single autoregressive Transformer, guiding the model to learn disentangled dynamics. As a result, FlowVLA produces coherent visual predictions and facilitates more efficient policy learning. Experiments on challenging robotics manipulation benchmarks demonstrate state-of-the-art performance with substantially improved sample efficiency, pointing toward a more principled foundation for world modeling. Project page: https://irpn-lab.github.io/FlowVLA/
@@ 2.1 Framework Overview
Figure 2: The FlowVLA Framework Architecture. Our model follows a two-stage paradigm. (Left) Stage 1: World Model Pre-training with Visual CoT. Input frames are encoded into appearance tokens (orange). The model then autoregressively predicts an interleaved sequence of motion tokens (cyan, representing flow) and future appearance tokens. This structured $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$ prediction forces the model to reason about dynamics before forecasting the future. (Right) Stage 2: Policy Fine-tuning. The pre-trained world model is adapted for control. Conditioned on a text instruction (green) and the current observation (orange), the model autoregressively predicts action tokens (blue) that are decoded into robot commands.
FlowVLA follows a two-stage training paradigm, consistent with state-of-the-art methods like UniVLA Wang et al. (2025) and WorldVLA Cen et al. (2025) to ensure a fair basis for comparison.
@@ 3.3 Analysis of World Modeling Capabilities (Q2)
Figure 3: Analysis of Physical Plausibility on the Bridge V2 Benchmark. This figure highlights common physical failures in the baseline model. In both examples, the baseline model (middle row) struggles to maintain physical consistency, leading to implausible outcomes such as a disappearing manipulator or erratic object behavior. In contrast, FlowVLA (bottom row), guided by its motion-first reasoning, produces stable and physically coherent predictions that accurately reflect the scene’s dynamics.
@@
Figure 4: Analysis of Semantic Alignment on the Bridge V2 Benchmark. This figure illustrates the baseline’s failure to align predictions with language instructions. While the predicted frames from baseline model (middle row) might appear visually plausible at a glance, the resulting motion does not correspond to the specified task (e.g., moving an object in the wrong direction). FlowVLA (bottom row) again demonstrates superior performance, correctly interpreting the command and generating a corresponding visual trajectory. This underscores that our Visual CoT not only improves physical realism but also enhances the model’s ability to ground language in action.
@@ 4.1 Vision-Language-Action (VLA) Models
The dominant paradigm for creating generalist robot agents is the Vision-Language-Action (VLA) model Zitkovich et al. (2023), Kim et al. (2024), Black et al. (2024), song2025acceleratingvisionlanguageactionmodelintegrated, song2025reconvla. These models extend large, pre-trained Vision-Language Models (VLMs) by fine-tuning them on extensive robotics datasets O’Neill et al. (2024). Architectures like RT-2 Zitkovich et al. (2023) and OpenVLA Kim et al. (2024) treat action generation as a sequence modeling problem, directly mapping visual and textual inputs to discretized action tokens. Other recent works have focused on improving the action representation itself, using techniques like diffusion policies Chi et al. (2023) or flow matching Black et al. (2024). While this end-to-end approach has demonstrated remarkable generalization, it often treats the environment’s physical dynamics as a “black box”. The policy is learned reactively, without an explicit, underlying model of how the world functions or evolves. FlowVLA diverges from this standard VLA formulation by prioritizing world understanding over immediate action generation. Its pre-training objective is not to learn a policy ($v_{t}\rightarrow a_{t}$), but to build a robust world model by learning the physical transition function of the environment ($v_{t}\rightarrow v_{t+1}$). This “dynamics-first” approach establishes a solid foundation of physical knowledge before it is adapted for downstream control.

### v2

@@ Abstract
Many Vision-Language-Action (VLA) models are built upon an internal world model trained via direct next-frame prediction ($v_{t}\rightarrow v_{t+1}$). This paradigm, however, presents a fundamental challenge: it conflates the task of predicting physical motion with that of rendering static appearance, forcing a single mechanism to handle both. This inherent coupling often leads to physically implausible forecasts and inefficient policy learning. To address this limitation, we introduce the Visual Chain of Thought (Visual CoT), a framework that disentangles these processes by compelling the model to first reason about motion dynamics before generating the future frame’s visual appearance. We instantiate this principle by proposing FlowVLA, an autoregressive Transformer that explicitly materializes this reasoning process as $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$”, where $f_{t}$ is an intermediate optical flow prediction. By forcing the model to first commit to a motion plan ($f_{t}$), FlowVLA learns disentangled dynamics, resulting in more coherent visual predictions and significantly more efficient policy learning. Experiments on challenging robotics manipulation benchmarks demonstrate that FlowVLA achieves state-of-the-art performance with substantially improved sample efficiency, pointing toward a more principled foundation for world modeling in VLAs. Project page: https://irpn-lab.github.io/FlowVLA/
@@ 2.1 Framework Overview
Figure 2: The FlowVLA Framework Architecture. Our model follows a two-stage paradigm. (Left) Stage 1: World Model Pre-training with Visual CoT. Input frames are encoded into appearance tokens (pink). The model then autoregressively predicts an interleaved sequence of motion tokens (blue, representing flow) and future appearance tokens. This structured $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$ prediction forces the model to reason about dynamics before forecasting the future. For conceptual clarity, the Image and Flow Tokenizers are visualized separately; in practice, they are the exact same module applied to both appearance and motion inputs. (Right) Stage 2: Policy Fine-tuning. The pre-trained world model is adapted for control. Conditioned on a text instruction (gray) and the current observation (pink), the model autoregressively predicts action tokens (green) that are decoded into robot actions.
As shown in Figure 2, FlowVLA follows a two-stage training paradigm, consistent with state-of-the-art methods like UniVLA Wang et al. (2025) and WorldVLA Cen et al. (2025) to ensure a fair basis for comparison.
@@ 3.3 Analysis of World Modeling Capabilities (Q2)
Figure 3: Analysis of Physical Plausibility on the Bridge V2 Dataset. This figure highlights common physical failures in the baseline model. In both examples, the baseline model (middle row) struggles to maintain physical consistency, leading to implausible outcomes such as a disappearing manipulator or erratic object behavior. In contrast, FlowVLA (bottom row), guided by its motion-first reasoning, produces stable and physically coherent predictions that accurately reflect the scene’s dynamics.
@@
Figure 4: Analysis of Semantic Alignment on the Bridge V2 Dataset. This figure illustrates the baseline’s failure to align predictions with language instructions. While the predicted frames from baseline model (middle row) might appear visually plausible at a glance, the resulting motion does not correspond to the specified task (e.g., moving an object in the wrong direction). FlowVLA (bottom row) again demonstrates superior performance, correctly interpreting the command and generating a corresponding visual trajectory. This underscores that our Visual CoT not only improves physical realism but also enhances the model’s ability to ground language in action.
@@ 4.1 Vision-Language-Action (VLA) Models
The dominant paradigm for creating generalist robot agents is the Vision-Language-Action (VLA) model Zitkovich et al. (2023), Kim et al. (2024), Black et al. (2024), Song et al. (2025b, c). These models extend large, pre-trained Vision-Language Models (VLMs) by fine-tuning them on extensive robotics datasets O’Neill et al. (2024). Architectures like RT-2 Zitkovich et al. (2023), CEED-VLA Song et al. (2025a) and OpenVLA Kim et al. (2024) treat action generation as a sequence modeling problem, directly mapping visual and textual inputs to discretized action tokens. Other recent works have focused on improving the action representation itself, using techniques like diffusion policies Chi et al. (2023) or flow matching Black et al. (2024). While this end-to-end approach has demonstrated remarkable generalization, it often treats the environment’s physical dynamics as a “black box”. The policy is learned reactively, without an explicit, underlying model of how the world functions or evolves. FlowVLA diverges from this standard VLA formulation by prioritizing world understanding over immediate action generation. Its pre-training objective is not to learn a policy ($v_{t}\rightarrow a_{t}$), but to build a robust world model by learning the physical transition function of the environment ($v_{t}\rightarrow v_{t+1}$). This “dynamics-first” approach establishes a solid foundation of physical knowledge before it is adapted for downstream control.